OpenStudy (anonymous):

use echelon method to solve the following system of three equations 4x-2y+3z=58 3x-2y-5z=-16 x-y+3z=3

OpenStudy (anonymous):

.isnt there a way to do this on graphing calc.

OpenStudy (kropot72):

Is the last equation in the question x - 7 + 3z = 31? Or should it be x -7y + 3z = 31?

OpenStudy (anonymous):

x-y+3z=31 sorry musta been a typo

OpenStudy (anonymous):

how do i determine what to multiply by